# Dept. of Informatics King’s College London London, UK

### v1

@@ 1 Introduction
However, more recently, the anthropocentric grip on understanding deception has weakened. Research on deception (and its detection) is expanding beyond human agents, to deceptive technologies, due to the current hybridisation of our societies. Hybrid societies are ‘self-organizing, collective systems, which are composed of different components, for example, natural and artificial parts (bio-hybrid) or human beings interacting with and through technical systems (socio-technical)’ Hamann et al. (2016). Nowadays, AI technologies play a crucial role in hybrid societies, but research in AI and deception has not progressed enough to allow us to understand and predict how advancements in the design of AI agents will impact hybrid societies.
A particular threat to the hybridisation of societies is the development of fully autonomous deceptive AI agents that will be able to form their own reasons and methods to perform deception, as well as out-think and outsmart humans and other AI agents Sarkadi (2021). By fully autonomous deceptive AI agent we mean neither the already existing human-scripted ‘mindless’ chatterbots which follow a pre-programmed script to deceive Mauldin (1994), nor the ‘clueless’ stochastic parrots Bender et al. (2021) which blurt out sentences without having any sense of their meaning in-context, but AI agents in the likes of the conceptual machines that trick the judges in the Imitation Game Turing (1950). In particular, we mean AI agents in the likes of the ones formally described by Cohen and Levesque (1985) or in the likes of Hamblin machines Staines (2018), which, in addition to being able to reason about what they communicate to others in various contexts and situations and how their communicative actions cause changes in the minds of others, also have deceptive motives.
@@
Research in multi-agent systems (MAS) aims to build models that integrate the social, behavioural and cognitive components of trust and deception. About 20 years ago, Castelfranchi and Tan (2001, 2002) emphasised that this is most important in hybrid human-agent interactions, raising the issue that in order for agents (human or artificial) to reason about the trustworthiness of their counterparts in different contexts, a theory of both trust and deception is necessary. Castelfranchi and Tan (2002) indicated multiple levels of trust such as trust in one’s agent and mediating agents, trust in the MAS environment and infrastructure, trust in potential partner (collaborator) agents, and finally trust in authorities. Their perspective is reinforced by the one presented by Falcone et al. (2001) on trust in cyber-societies, which is becoming more relevant given the increasing number and complexity of hybrid interactions between humans and artificial agents. So far, most work has focused on modelling various aspects of trust, independent of deception, as can be observed by comparing the trend in Fig.1(a) to the one in Fig.1(b). Hence, the development of deceptive machines for the purpose of building a holistic socio-cognitive theory has been scarce, and when it did happen, the focus was mostly on particular and very narrow aspects of deception.
@@
Why is this a problem? Because due to its emergence in hybrid societies, deception can lead to a Tragedy of The Digital Commons, where AI agents maliciously exploit and pollute information shared with humans and other AI agents Greco and Floridi (2004); Sarkadi et al. (2021a). If we want to understand its conditions, consequences, when, why, how, to whom it happens and who is responsible for it, then most of deception’s components and forms cannot be treated independently.
@@
In the real world, the area of Intelligence Analysis tries to understand complex phenomena holistically - in various scenarios and contexts, and weighs the risks and impacts of these phenomena. One of the most difficult of these phenomena to explain is considered to be the one of deception. Hence, to understand it holistically, the Intelligence community has developed rigorous methodologies for performing the process of inference to the best explanation (IBE) Heuer (1999). But how do analysts actually explain and make sense of complex and potentially deceptive agent interactions? According to Ferris (1989), intelligence analysts and historians of deception try to understand two processes, namely what causes deception and what prevents deception from being caused. Causation and causation prevention of deception also need to be understood in context and as part of an overarching narrative. The main methods for understanding deception in Intelligence Analysis come in the form of structured analytic techniques (SATs - not to be confused with SAT solvers in Computer Science!), most notably the Analysis of Competing Hypotheses (ACH) and its derivations that are used by intelligence analysts to detect deception and reduce cognitive load and bias Heuer (1999). However, despite being designed to exhaust all possible scenarios of deception and deception causation prevention in complex cases, these methods have proved to be rather tedious to perform by analysts - they are cognitively demanding Van Gelder (2008); Pope et al. (2006).
@@ 2.1 Philosophical Perspectives
The following, epistemically exhaustive definition of deception is provided by Chisholm and Feehan (1977), and by Mahon (2016):
@@
If we are careful, we can observe that the definition above merely enumerates the modalities of deception. That is, it describes a category of dishonest communicative behaviours, e.g. to palter, to pander, to lie by omission, to lie by commission, telling half-truths, etc. that can be used to deceive. While an enumeration or a taxonomy of all the ways in which a phenomenon may be instantiated can be useful to understand the phenomenon of deception, according to some researchers in Communication Theory it is by no means a good or proper definition of the phenomenon in itself Levine (2019). WhatLevine (2019) suggests is that a functional definition of deception would be more appropriate than merely laying down a behavioural taxonomy. A functional perspective on deception is also endorsed in Philosophy by Artiga and Paternotte (2018). Taking their suggestions into account, a functional definition that we adopt to define deception and to refer to deception throughout the paper, without reducing the scope of the discussion to specific types of communicative social behaviour that might be used for deception, is the following:
@@ 2.2 Socio-Cognitive Perspectives
This simplistic view of deception, especially the over-reliance on non-verbal cues as micro-expressions, had been extremely popular for some time due to the work of Ekman and Friesen (1969). Today, however, there is relatively strong consensus in human deception research that cue-based approaches are very limited and can help detect deception only slightly better than chance Levine (2015). Notably, the work of DePaulo et al. (1996, 2003); Bond Jr and DePaulo (2008), has shed some light on how the over-reliance on specific cues is problematic and that these do not improve deception detection. This makes cue-based deception research highly susceptible to cognitive biases such as the truth bias or the confirmation bias (a.k.a. the Othello Error) Bond Jr and Fahey (1987).
@@
Alternatively, non-cue based approaches, instead of trying to solely focus on deception detection, they aim to enrich the understanding of various factors and components that underlie social interaction. The following non-cue based theories are actually the ones that can provide with a truly socio-cognitive understanding of deception: Information Manipulation Theory 2 (IMT2) proposed by McCornack et al. (2014) and Truth-Default Theory (TDT) proposed by Levine (2014). IMT2 aims to understand the cognitive processes that are responsible with speech production in deceptive interactions. One of the arguments proposed by IMT2 is that there is no difference w.r.t. cognitive load when humans lie or tell the truth. TDT differs from IMT2, by focusing on the contextual aspects of deception. TDT proposes that there is a general truth-default state in which humans are susceptible to deception and from which they can be triggered out of depending on the informational context. In addition to IMT2 and TDT, there is Interpersonal Deception Theory (IDT) proposed by Burgoon et al. (1996), which can provide us with a rich socio-cognitive perspective. While IDT is cue-based to some extent, it is not over-reliant on cues. This theory is very different from standard cue-based approaches and its strength is that it aims to give an integrated perspective on how cues influence inter-personal interactions between human agents. In one way, IDT represents deceptive interactions as a multi-agent system that is governed by rules of message transfer.
@@
There is a common component throughout all of the non-cue based perspectives on deception, namely the ability to form, simulate, and reason about the minds of others, which is known in the scientific literature as the ability to use Theory-of-Mind (ToM). Different ‘flavours’ of ToM have been proposed to explain this phenomenon. One flavour is Theory-Theory of Mind (TT). TT represents the type of ToM where agents already have an understanding to some extent of the other agents’ minds. TT can be, for instance, a set or a knowledge base that consists of beliefs of others’ mental attitudes. Another flavour is known as Simulation Theory of Mind (ST). ST represents a process for using the already known set of others’ mental attitudes to simulate their minds in order to see what other attitudes they can form or to predict their behaviour. Both types of ToM have emerged from the study of children in developmental psychology Gopnik and Wellman (1992). Gopnik and Wellman describe how the two competing flavours aim to explain how children predict others’ behaviour, e.g. whether they use language associated to TT such as ‘beliefs’ and ‘desires’ - theoretical language constructs that form a conceptual representation of another’s mind, or whether they use phenomenal constructs associated to ST by using their own minds to ‘run’ a simulation of inputs and generate outputs in the form of an ‘experience’. In TT, the child heavily relies on a conceptual constructs and language, whereas in ST egocentrism plays a crucial role, as the child heavily relies on projecting one’s own mental states onto another.
@@
Even if there has been an ongoing debate namely between the proponents of TT and the proponents of ST regarding how humans develop and use ToM, these perspectives seem to be merging into a hybrid perspective on ToM Goldman et al. (2012). While TT represents the type of ToM where agents already have an understanding to some extent of the other agents’ minds, e.g., a set or a knowledge base that consists of beliefs of others’ mental attitudes, ST represents a process for using the already known set of others’ mental attitudes to simulate their minds in order to see what other attitudes they can form or to predict their behaviour. TT and ST can be merged into hybrid ToM, which can be thought of as a ‘high-level’ ToM in the form of a simulation setup in which an agent takes its target’s beliefs, desires, and intentions as premises for generating a practical reasoning process in the way its target (the other agent who’s mind is being mentalised) would Goldman et al. (2006, 2012).
@@
Some of the other socio-cognitive components of deception have been described by IMT2 to consider how information or knowledge is managed inside the deceivers’ cognitive apparatus. This cognitive apparatus in humans represents some sort of cognitive agent architecture that is able to run various processes before, after, or during communicative interactions. IMT2 describes it as ‘a speech production system involving parallel-distributed-processing guided by efficiency, memory, and means-ends reasoning; and this production process involves a rapid-fire series of cognitive cycles (involving distinct modules united by a conscious work-space), and modification of incrementally-constructed discourse during the turn-at-talk in response to dynamic current-state/end-state discrepancies’. IMT2 identifies two main cognitive processes derived from speech-act theory of cognition that are responsible for information selection and dissemination. The first process is called Pars Pro Toto, meaning the parts for the whole, which a sender agent uses to select the information that, given the context in which the sender and the receiver find themselves, will convey the sender’s intended meaning of the message in the mind of the receiver. The second process is called Totum Ex Parte, meaning the whole from the parts, which a receiver uses to infer the intended meaning of the message from a sender given the context in which the receiver and the sender find themselves in. According to IMT2, deceivers and deception detectors engage in these two processes when they interact. Fascinatingly, IMT2 has been inspired by work at the intersection of AI and pragmatics Newell et al. (1972); Hovy (1990), hence its adoption of specialised terms from AI when describing the speech production systems for humans. Unsurprisingly to people working in AI, IMT2 seems highly tuned to the idea of intelligent socially-aware AI agents Cohen and Levesque (1985).
@@
Furthermore, TDT Levine (2019) describes the contextual level of deception, which informs us that there is more to deceptive interactions than social parameters like cognitive load and leakage when it comes to interpreting the information received or transmitted. According to TDT, what is communicated (informational content) in a given context should be given more importance than non-verbal behaviour that may or may not happen to be correlated with the communicated content. Therefore, based on TDT, the contextual knowledge and information available to the agents that interact with each other should be a crucial component of computational deception. The way this information is used can determine whether deceptive attempts are successful or not. The same goes for attempts at detecting deception, such as persuading an agent to reveal deceptive motives. This information should represent knowledge about what is said or not, when and where it is said (or not), to whom (and not), as well as how it is or not interpreted by someone. This type of information can trigger agents into or out of the truth-default mental state, which makes individuals spend less or more cognitive resources on what is being communicated, e.g. makes them more or less suspicious to others depending on the context.
Even though IMT2 and IDT do not mention ToM as an explicit component of deception, ToM seems to be assumed implicitly. For instance, as a ‘high-level simulation theory’, hybrid ToM blends TT and ST into practical reasoning Goldman et al. (2006, 2012), which is a process very similar, if not identical, to the processes of Pars Pro Toto and Totum Ex Parte described by IMT2. Now, the only theory of deception out of the three described here, that explicitly mentions ToM is TDT Levine (2019), where deception would be a special case of ToM. This seems to be a well-founded consideration given that humans develop the ability to form and use ToM in early childhood. Another strong argument for the tight link between human ToM and deception is the use of false-belief experiments to track the formation of ToM in children Gopnik and Wellman (2012). Based on this, we can look at the definition of deception (see Section 2.1) and safely say that intentional exploitation of false belief formation is nothing other than deception, confirming Levine’s point that deception is a particular case of ToM in use.
@@ 3.1 Multi-Agent Frameworks vs AI Reductionism
There exists a methodological, or better said a paradigmatic limitation to building holistic frameworks for modelling deception in AI. This limitation comes from the fact that AI approaches are usually highly reductionistic, namely that real world factors are abstracted away for the purpose of “elegance” and simplicity of AI models. In the history of science and philosophy, reductionism has been extensively debated for hundreds of years Andersen (2001). In modern science, Marr (2010) started discussing the limitations of reducing cognitive functions to physical properties of neurons, arguing that explanations about cognitive functions cannot solely rely on neuro-chemical activity Peebles and Cooper (2015); Bickle (2015), but that in order to be meaningfully understood they also require references to behaviour as part of an overarching multi-level analysis Marr (2010); Krakauer et al. (2017).
A similar argument can be made about research in AI. That is if we want to understand and explain a socio-cognitive phenomenon such as deception, for which other areas of scientific and humanities research have identified various components that interact with each other to give rise to a complex multi-layered process inside complex multi-layered systems that are hybrid societies, then we cannot reduce these types of processes to a single function. Historically, AI has been all about breaking down human intelligence and modelling it computationally as functions at a very abstract level Boden (1996). This is understandable for investigating the properties of some processes in isolation. But when a process consists of a multitude of these functions working sequentially and in parallel at multiple levels of abstraction - from the internal cognitive reasoning mechanisms to the interaction mechanisms responsible for communication to the emergent large-scale behaviour of agents that are interacting - then just solely focusing on the properties of a single computational function, independently of other components, seems to miss the point of modelling the entire phenomenon itself [^1] The same argument applies to why definitions that explain the entire phenomenon of deception are better suited than others which only define a particular type of deceptive behaviour.. Aside from representing the functions themselves, there is the problem of a causality between elements at various layers of abstraction, e.g. the problem of how functions relate to each other. In turn, this means that AI models need to account for different kinds of architectures. This is why the area of Distributed AI, and later ABM and MAS, emerged to be able to study systems of agents. However, even in these areas, most of the work has lost focus of socio-cognitive modelling and has turned to benchmarking and optimisation of groups of artificial agents Dignum and Dignum (2020).
@@
One method in MAS to limit the negative effects of reductionism is to model and integrate layers of abstraction, which are the foundations of a MAS framework Boissier et al. (2013). Their integration accounts for a rich multi-dimensional world instantiated through computational representations of environments, agents, and organisations. Social interactions between agents are complex and some of the properties of these interactions and the relations between these properties can only be described by taking into account these different layers. As an illustrative example from MAS engineering, the JaCaMo framework  Boissier et al. (2013) integrates the three layers of abstraction respectively through the Jason agent oriented programming language (agent layer), the CArtAgO common artifact infrastructure for agents open environments (environment layer), and the Moise organisation oriented programming framework (organisation layer). Various MAS models can be then represented and implemented using such a framework to explore agent systems at the level of interaction, considering how agents interact with the environment and each other as part of various organisations that follow different norms etc.
Another way to deal with reductionism in AI is to account for multiple models and levels of abstraction through a combination of MAS and ABM approaches, representations, and techniques, formal or computational, that can be used to describe the properties of interactions between agents. MAS approaches focus on modelling the interaction level, while ABM approaches focus on modelling the descriptive level Dignum and Dignum (2020). An example of a MAS model that focuses on interaction between agents is a computational model based on speech-acts. Such a model describes and/or predicts both the cognitive processes and the behaviour used to produce changes in the belief bases of other agents through speech-production. This type of model, for instance, can be used to track the information exchanged during a dialogue, as well as the reasoning that the agents perform during the dialogue. On the other hand, an ABM model that focuses on describing emergent behaviour does not need to represent every cognitive component responsible for producing a speech-act, it only needs to describe an overall state of a system. Thus, it can be efficiently used to represent the costs and benefits of communication and compute states of the system at different moments through software simulation. For instance, it can check if an equilibrium is ever reached for exchanging information given a set of social parameters.

### v2

@@ 1 Introduction
However, more recently, the anthropocentric grip on understanding deception has weakened. Research on deception (and its detection) is expanding beyond human agents, to deceptive technologies, due to the current hybridisation of our societies. Hybrid societies are ‘self-organizing, collective systems, which are composed of different components, for example, natural and artificial parts (bio-hybrid) or human beings interacting with and through technical systems (socio-technical)’ (Hamann et al., 2016). Nowadays, AI technologies play a crucial role in hybrid societies, but research in AI and deception has not progressed enough to allow us to understand and predict how advancements in the design of AI agents will impact hybrid societies.
A particular threat to the hybridisation of societies is the development of fully autonomous deceptive AI agents that will be able to form their own reasons and methods to perform deception, as well as out-think and outsmart humans and other AI agents (Sarkadi, 2021). By fully autonomous deceptive AI agent we mean neither the already existing human-scripted ‘mindless’ chatterbots which follow a pre-programmed script to deceive (Mauldin, 1994), nor the ‘clueless’ stochastic parrots (Bender et al., 2021) which blurt out sentences without having any sense of their meaning in-context, but AI agents in the likes of the conceptual machines that trick the judges in the Imitation Game (Turing, 1950). In particular, we mean AI agents in the likes of the ones formally described by (Cohen and Levesque, 1985) or in the likes of Hamblin machines (Staines, 2018), which, in addition to being able to reason about what they communicate to others in various contexts and situations and how their communicative actions cause changes in the minds of others, also have deceptive motives.
@@
Research in multi-agent systems (MAS) aims to build models that integrate the social, behavioural and cognitive components of trust and deception. About 20 years ago, (Castelfranchi and Tan, 2001, 2002) emphasised that this is most important in hybrid human-agent interactions, raising the issue that in order for agents (human or artificial) to reason about the trustworthiness of their counterparts in different contexts, a theory of both trust and deception is necessary. (Castelfranchi and Tan, 2002) indicated multiple levels of trust such as trust in one’s agent and mediating agents, trust in the MAS environment and infrastructure, trust in potential partner (collaborator) agents, and finally trust in authorities. Their perspective is reinforced by the one presented by (Falcone et al., 2001) on trust in cyber-societies, which is becoming more relevant given the increasing number and complexity of hybrid interactions between humans and artificial agents. So far, most work has focused on modelling various aspects of trust, independent of deception, as can be observed by comparing the trend in Fig.1(a) to the one in Fig.1(b). Hence, the development of deceptive machines for the purpose of building a holistic socio-cognitive theory has been scarce, and when it did happen, the focus was mostly on particular and very narrow aspects of deception.
@@
Why is this a problem? Because due to its emergence in hybrid societies, deception can lead to a Tragedy of The Digital Commons, where AI agents maliciously exploit and pollute information shared with humans and other AI agents (Greco and Floridi, 2004; Sarkadi et al., 2021a). If we want to understand its conditions, consequences, when, why, how, to whom it happens and who is responsible for it, then most of deception’s components and forms cannot be treated independently.
@@
In the real world, the area of Intelligence Analysis tries to understand complex phenomena holistically - in various scenarios and contexts, and weighs the risks and impacts of these phenomena. One of the most difficult of these phenomena to explain is considered to be the one of deception. Hence, to understand it holistically, the Intelligence community has developed rigorous methodologies for performing the process of inference to the best explanation (IBE) (Heuer, 1999). But how do analysts actually explain and make sense of complex and potentially deceptive agent interactions? According to (Ferris, 1989), intelligence analysts and historians of deception try to understand two processes, namely what causes deception and what prevents deception from being caused. Causation and causation prevention of deception also need to be understood in context and as part of an overarching narrative. The main methods for understanding deception in Intelligence Analysis come in the form of structured analytic techniques (SATs - not to be confused with SAT solvers in Computer Science!), most notably the Analysis of Competing Hypotheses (ACH) and its derivations that are used by intelligence analysts to detect deception and reduce cognitive load and bias (Heuer, 1999). However, despite being designed to exhaust all possible scenarios of deception and deception causation prevention in complex cases, these methods have proved to be rather tedious to perform by analysts - they are cognitively demanding (Van Gelder, 2008; Pope et al., 2006).
@@ 2.1 Philosophical Perspectives
The following, epistemically exhaustive definition of deception is provided by (Chisholm and Feehan, 1977), and by (Mahon, 2016):
@@
If we are careful, we can observe that the definition above merely enumerates the modalities of deception. That is, it describes a category of dishonest communicative behaviours, e.g. to palter, to pander, to lie by omission, to lie by commission, telling half-truths, etc. that can be used to deceive. While an enumeration or a taxonomy of all the ways in which a phenomenon may be instantiated can be useful to understand the phenomenon of deception, according to some researchers in Communication Theory it is by no means a good or proper definition of the phenomenon in itself (Levine, 2019). What(Levine, 2019) suggests is that a functional definition of deception would be more appropriate than merely laying down a behavioural taxonomy. A functional perspective on deception is also endorsed in Philosophy by (Artiga and Paternotte, 2018). Taking their suggestions into account, a functional definition that we adopt to define deception and to refer to deception throughout the paper, without reducing the scope of the discussion to specific types of communicative social behaviour that might be used for deception, is the following:
@@ 2.2 Socio-Cognitive Perspectives
This simplistic view of deception, especially the over-reliance on non-verbal cues as micro-expressions, had been extremely popular for some time due to the work of (Ekman and Friesen, 1969). Today, however, there is relatively strong consensus in human deception research that cue-based approaches are very limited and can help detect deception only slightly better than chance (Levine, 2015). Notably, the work of (DePaulo et al., 1996, 2003; Bond Jr and DePaulo, 2008), has shed some light on how the over-reliance on specific cues is problematic and that these do not improve deception detection. This makes cue-based deception research highly susceptible to cognitive biases such as the truth bias or the confirmation bias (a.k.a. the Othello Error) (Bond Jr and Fahey, 1987).
@@
Alternatively, non-cue based approaches, instead of trying to solely focus on deception detection, they aim to enrich the understanding of various factors and components that underlie social interaction. The following non-cue based theories are actually the ones that can provide with a truly socio-cognitive understanding of deception: Information Manipulation Theory 2 (IMT2) proposed by (McCornack et al., 2014) and Truth-Default Theory (TDT) proposed by (Levine, 2014). IMT2 aims to understand the cognitive processes that are responsible with speech production in deceptive interactions. One of the arguments proposed by IMT2 is that there is no difference w.r.t. cognitive load when humans lie or tell the truth. TDT differs from IMT2, by focusing on the contextual aspects of deception. TDT proposes that there is a general truth-default state in which humans are susceptible to deception and from which they can be triggered out of depending on the informational context. In addition to IMT2 and TDT, there is Interpersonal Deception Theory (IDT) proposed by (Burgoon et al., 1996), which can provide us with a rich socio-cognitive perspective. While IDT is cue-based to some extent, it is not over-reliant on cues. This theory is very different from standard cue-based approaches and its strength is that it aims to give an integrated perspective on how cues influence inter-personal interactions between human agents. In one way, IDT represents deceptive interactions as a multi-agent system that is governed by rules of message transfer.
@@
There is a common component throughout all of the non-cue based perspectives on deception, namely the ability to form, simulate, and reason about the minds of others, which is known in the scientific literature as the ability to use Theory-of-Mind (ToM). Different ‘flavours’ of ToM have been proposed to explain this phenomenon. One flavour is Theory-Theory of Mind (TT). TT represents the type of ToM where agents already have an understanding to some extent of the other agents’ minds. TT can be, for instance, a set or a knowledge base that consists of beliefs of others’ mental attitudes. Another flavour is known as Simulation Theory of Mind (ST). ST represents a process for using the already known set of others’ mental attitudes to simulate their minds in order to see what other attitudes they can form or to predict their behaviour. Both types of ToM have emerged from the study of children in developmental psychology (Gopnik and Wellman, 1992). Gopnik and Wellman describe how the two competing flavours aim to explain how children predict others’ behaviour, e.g. whether they use language associated to TT such as ‘beliefs’ and ‘desires’ - theoretical language constructs that form a conceptual representation of another’s mind, or whether they use phenomenal constructs associated to ST by using their own minds to ‘run’ a simulation of inputs and generate outputs in the form of an ‘experience’. In TT, the child heavily relies on a conceptual constructs and language, whereas in ST egocentrism plays a crucial role, as the child heavily relies on projecting one’s own mental states onto another.
@@
Even if there has been an ongoing debate namely between the proponents of TT and the proponents of ST regarding how humans develop and use ToM, these perspectives seem to be merging into a hybrid perspective on ToM (Goldman et al., 2012). While TT represents the type of ToM where agents already have an understanding to some extent of the other agents’ minds, e.g., a set or a knowledge base that consists of beliefs of others’ mental attitudes, ST represents a process for using the already known set of others’ mental attitudes to simulate their minds in order to see what other attitudes they can form or to predict their behaviour. TT and ST can be merged into hybrid ToM, which can be thought of as a ‘high-level’ ToM in the form of a simulation setup in which an agent takes its target’s beliefs, desires, and intentions as premises for generating a practical reasoning process in the way its target (the other agent who’s mind is being mentalised) would (Goldman et al., 2006, 2012).
@@
Some of the other socio-cognitive components of deception have been described by IMT2 to consider how information or knowledge is managed inside the deceivers’ cognitive apparatus. This cognitive apparatus in humans represents some sort of cognitive agent architecture that is able to run various processes before, after, or during communicative interactions. IMT2 describes it as ‘a speech production system involving parallel-distributed-processing guided by efficiency, memory, and means-ends reasoning; and this production process involves a rapid-fire series of cognitive cycles (involving distinct modules united by a conscious work-space), and modification of incrementally-constructed discourse during the turn-at-talk in response to dynamic current-state/end-state discrepancies’. IMT2 identifies two main cognitive processes derived from speech-act theory of cognition that are responsible for information selection and dissemination. The first process is called Pars Pro Toto, meaning the parts for the whole, which a sender agent uses to select the information that, given the context in which the sender and the receiver find themselves, will convey the sender’s intended meaning of the message in the mind of the receiver. The second process is called Totum Ex Parte, meaning the whole from the parts, which a receiver uses to infer the intended meaning of the message from a sender given the context in which the receiver and the sender find themselves in. According to IMT2, deceivers and deception detectors engage in these two processes when they interact. Fascinatingly, IMT2 has been inspired by work at the intersection of AI and pragmatics (Newell et al., 1972; Hovy, 1990), hence its adoption of specialised terms from AI when describing the speech production systems for humans. Unsurprisingly to people working in AI, IMT2 seems highly tuned to the idea of intelligent socially-aware AI agents (Cohen and Levesque, 1985).
@@
Furthermore, TDT (Levine, 2019) describes the contextual level of deception, which informs us that there is more to deceptive interactions than social parameters like cognitive load and leakage when it comes to interpreting the information received or transmitted. According to TDT, what is communicated (informational content) in a given context should be given more importance than non-verbal behaviour that may or may not happen to be correlated with the communicated content. Therefore, based on TDT, the contextual knowledge and information available to the agents that interact with each other should be a crucial component of computational deception. The way this information is used can determine whether deceptive attempts are successful or not. The same goes for attempts at detecting deception, such as persuading an agent to reveal deceptive motives. This information should represent knowledge about what is said or not, when and where it is said (or not), to whom (and not), as well as how it is or not interpreted by someone. This type of information can trigger agents into or out of the truth-default mental state, which makes individuals spend less or more cognitive resources on what is being communicated, e.g. makes them more or less suspicious to others depending on the context.
Even though IMT2 and IDT do not mention ToM as an explicit component of deception, ToM seems to be assumed implicitly. For instance, as a ‘high-level simulation theory’, hybrid ToM blends TT and ST into practical reasoning (Goldman et al., 2006, 2012), which is a process very similar, if not identical, to the processes of Pars Pro Toto and Totum Ex Parte described by IMT2. Now, the only theory of deception out of the three described here, that explicitly mentions ToM is TDT (Levine, 2019), where deception would be a special case of ToM. This seems to be a well-founded consideration given that humans develop the ability to form and use ToM in early childhood. Another strong argument for the tight link between human ToM and deception is the use of false-belief experiments to track the formation of ToM in children (Gopnik and Wellman, 2012). Based on this, we can look at the definition of deception (see Section 2.1) and safely say that intentional exploitation of false belief formation is nothing other than deception, confirming Levine’s point that deception is a particular case of ToM in use.
@@ 3.1 Multi-Agent Frameworks vs AI Reductionism
There exists a methodological, or better said a paradigmatic limitation to building holistic frameworks for modelling deception in AI. This limitation comes from the fact that AI approaches are usually highly reductionistic, namely that real world factors are abstracted away for the purpose of “elegance” and simplicity of AI models. In the history of science and philosophy, reductionism has been extensively debated for hundreds of years (Andersen, 2001). In modern science, (Marr, 2010) started discussing the limitations of reducing cognitive functions to physical properties of neurons, arguing that explanations about cognitive functions cannot solely rely on neuro-chemical activity (Peebles and Cooper, 2015; Bickle, 2015), but that in order to be meaningfully understood they also require references to behaviour as part of an overarching multi-level analysis (Marr, 2010; Krakauer et al., 2017).
A similar argument can be made about research in AI. That is if we want to understand and explain a socio-cognitive phenomenon such as deception, for which other areas of scientific and humanities research have identified various components that interact with each other to give rise to a complex multi-layered process inside complex multi-layered systems that are hybrid societies, then we cannot reduce these types of processes to a single function. Historically, AI has been all about breaking down human intelligence and modelling it computationally as functions at a very abstract level (Boden, 1996). This is understandable for investigating the properties of some processes in isolation. But when a process consists of a multitude of these functions working sequentially and in parallel at multiple levels of abstraction - from the internal cognitive reasoning mechanisms to the interaction mechanisms responsible for communication to the emergent large-scale behaviour of agents that are interacting - then just solely focusing on the properties of a single computational function, independently of other components, seems to miss the point of modelling the entire phenomenon itself [^1] The same argument applies to why definitions that explain the entire phenomenon of deception are better suited than others which only define a particular type of deceptive behaviour.. Aside from representing the functions themselves, there is the problem of a causality between elements at various layers of abstraction, e.g. the problem of how functions relate to each other. In turn, this means that AI models need to account for different kinds of architectures. This is why the area of Distributed AI, and later ABM and MAS, emerged to be able to study systems of agents. However, even in these areas, most of the work has lost focus of socio-cognitive modelling and has turned to benchmarking and optimisation of groups of artificial agents (Dignum and Dignum, 2020).
@@
One method in MAS to limit the negative effects of reductionism is to model and integrate layers of abstraction, which are the foundations of a MAS framework (Boissier et al., 2013). Their integration accounts for a rich multi-dimensional world instantiated through computational representations of environments, agents, and organisations. Social interactions between agents are complex and some of the properties of these interactions and the relations between these properties can only be described by taking into account these different layers. As an illustrative example from MAS engineering, the JaCaMo framework  (Boissier et al., 2013) integrates the three layers of abstraction respectively through the Jason agent oriented programming language (agent layer), the CArtAgO common artifact infrastructure for agents open environments (environment layer), and the Moise organisation oriented programming framework (organisation layer). Various MAS models can be then represented and implemented using such a framework to explore agent systems at the level of interaction, considering how agents interact with the environment and each other as part of various organisations that follow different norms etc.
Another way to deal with reductionism in AI is to account for multiple models and levels of abstraction through a combination of MAS and ABM approaches, representations, and techniques, formal or computational, that can be used to describe the properties of interactions between agents. MAS approaches focus on modelling the interaction level, while ABM approaches focus on modelling the descriptive level (Dignum and Dignum, 2020). An example of a MAS model that focuses on interaction between agents is a computational model based on speech-acts. Such a model describes and/or predicts both the cognitive processes and the behaviour used to produce changes in the belief bases of other agents through speech-production. This type of model, for instance, can be used to track the information exchanged during a dialogue, as well as the reasoning that the agents perform during the dialogue. On the other hand, an ABM model that focuses on describing emergent behaviour does not need to represent every cognitive component responsible for producing a speech-act, it only needs to describe an overall state of a system. Thus, it can be efficiently used to represent the costs and benefits of communication and compute states of the system at different moments through software simulation. For instance, it can check if an equilibrium is ever reached for exchanging information given a set of social parameters.